C3H8 + 5 O2 ---- > 3 CO2 + 4 H2O
Calculate the mass of CO2 produced when 6.5 g of propane is reacted with 14.2 g of O2.
1
Expert's answer
2020-06-29T07:52:49-0400

m(CO2)=m(O2)*3Mr(CO2)/5Mr(O2)=14.2*3*44/5*32=11.715 g;

Answer: 11.715 g